OpenStudy (anonymous):

find potential energy of a 96 N sack of grain hoisted to a storage room 52 m above the ground floor of a grain elevator.

OpenStudy (abhisar):

Potential energy stored on an object = Work done on it \(\sf \therefore\) Potential energy = Force \(\times\) distance = Mass \(\times\)g\(\times\)Height

OpenStudy (abhisar):

Getting it?

OpenStudy (anonymous):

yes thank you very much

OpenStudy (abhisar):

Welcome :)